# First-Spammed, First-Served: MEV Extraction on Fast-Finality Blockchains

### v2

@@ 1 Introduction
Section 2 provides an overview of roll-ups, Layer-2 (L2) blockchain mechanisms, and their implications for MEV dynamics. Section 3 builds the optimal MEV strategy on fast-finality blockchain for a MEV arbitrageur. Subsequently, sections 4LABEL:sec:methodology and 5 present an empirical analysis of reverted transactions on L2s and explain their link to MEV extraction. Finally, Sections 6 and 7 include the discussion and conclusions.
@@ 4 Data Collection and Methodology
Data Collection: A key challenge in our empirical analysis is that DeFi smart contracts do not emit event logs for reverted transactions. To overcome this, we analyze transaction traces directly. For each reverted transaction, we construct a trace-based execution graph and search it to identify DEX or router interactions, targeted liquidity pools (Uniswap v2/v3), and token addresses (Uniswap v4). Although all analyzed rollups are EVM-compatible, each liquidity pool, AMM DEX, router, and token contract has a different address across L2s, requiring chain-specific mapping. To support this, we build a labeling library for each rollup using event logs from successfully executed swaps, which we then use to match and annotate nodes in the trace graphs of reverted transactions. In this study, we analyze both successful and reverted transactions on leading EVM-compatible rollups: Arbitrum, Optimism, Base, ZKsync, and Unichain. The dataset spans from the Dencun upgrade (March 2023) to the present and is sourced from full archive nodes. A detailed overview of the analyzed rollups is provided in Table 1. Notably, all analyzed rollups have private mempools operated by centralized sequencers with block times ranging between 0.2s and 2s (compared to 12s in Ethereum). The default order flow is First-Come First-Served (FCFS) with Priority Fee Auctions (PFA). On April 17, 2025, Arbitrum introduced the TimeBoost mechanism (explained further).
@@
Methodology: A key challenge in our empirical analysis is that DeFi smart contracts do not emit event logs for reverted transactions. To overcome this, we analyze transaction traces directly. For each reverted transaction, we construct a trace-based execution graph and search it to identify DEX or router interactions, targeted liquidity pools (Uniswap v2/v3), and token addresses (Uniswap v4). Although all analyzed rollups are EVM-compatible, each liquidity pool, AMM DEX, router, and token contract has a different address across L2s, requiring chain-specific mapping. To support this, we build a labeling library for each rollup using event logs from successfully executed swaps, which we then use to match and annotate nodes in the trace graphs of reverted transactions.
Successful transactions: We extract swap details from Uniswap v2, v3, and v4 by parsing their emitted event logs.
Reverted transactions: Since reverted transactions do not emit logs, we use on transaction traces to analyze their behavior.
@@
Bots Identification: We identify bots based on the from address rather than to, as bots can be called from multiple sources. The further logic follows. If the address has a known label or owner name, it’s likely not a bot, e.g., labeled wallets from Etherscan, centralized exchanges, DAOs, etc. (owa.owner-key IS NOT NULL). If the address is a contract address (i.e., contains bytecode on-chain), it’s likely an automated agent, such as a bot or smart contract trader (cm.contract-address IS NOT NULL).
@@ 5 Empirical Analysis
In the subsequent sections, we analyze the reverted transactions with respect to their use of priority fee auctions and their position within blocks. We classify reverted transactions into swap transactions and examine their key characteristics, including the targeted decentralized exchange (DEX) and the token pair pool.
@@ 6 Discussion
The Dencun Effect
The Dencun upgrade introduced a significant shift in L2 gas economics. Pre-upgrade, the relatively high cost (often \$0.5–\$1 per transaction) disincentivized failed spam-based arbitrage. Post-Dencun, with average gas costs dropping below \$0.01, the marginal cost of failure became negligible. This cost asymmetry makes spam-based arbitrage economically rational post-upgrade and explains the observed spike in revert rates.
@@
Incentive Design and New Sequencing Mechanisms
The current status quo—FCFS ordering on centralized, private mempools—favors low-latency spamming and limits the scope for efficient, market-based prioritization. Several mechanisms have been proposed to address this:
@@
Underreliance on Priority Fee Auctions
A notable empirical observation is the limited use of priority fee auctions (PFA) by MEV searchers on rollups. Despite the theoretical promise of PFAs to prioritize transaction inclusion, we find that searchers often prefer to duplicate their transactions—submitting the same swap multiple times—rather than bid for priority.
@@ Appendix 0.A Gas Fees of Reverted Transactions
  • L2 Execution Fee: The cost of executing the transaction on the rollup’s virtual machine.
  • L1 Data Availability Fee: The cost of posting transaction calldata to Ethereum L1, which ensures data availability and security.

### v3

@@ 1 Introduction
Section 2 provides an overview of roll-ups, Layer-2 (L2) blockchain mechanisms, and their implications for MEV dynamics. Section 3 builds the optimal MEV strategy on fast-finality blockchain for a MEV arbitrageur. Subsequently, sections 4 and 5 present an empirical analysis of reverted transactions on L2s and explain their link to MEV extraction. Finally, Sections 6 and 7 include the discussion and conclusions.
@@ 4 Data Collection and Methodology
Data Collection. A key challenge in our empirical analysis is that DeFi smart contracts do not emit event logs for reverted transactions. To overcome this, we analyze transaction traces directly. For each reverted transaction, we construct a trace-based execution graph and search it to identify DEX or router interactions, targeted liquidity pools (Uniswap v2/v3), and token addresses (Uniswap v4). Although all analyzed rollups are EVM-compatible, each liquidity pool, AMM DEX, router, and token contract has a different address across L2s, requiring chain-specific mapping. To support this, we build a labeling library for each rollup using event logs from successfully executed swaps, which we then use to match and annotate nodes in the trace graphs of reverted transactions. In this study, we analyze both successful and reverted transactions on leading EVM-compatible rollups: Arbitrum, Optimism, Base, ZKsync, and Unichain. The dataset spans from the Dencun upgrade (March 2023) to the present and is sourced from full archive nodes. A detailed overview of the analyzed rollups is provided in Table 1. Notably, all analyzed rollups have private mempools operated by centralized sequencers with block times ranging between 0.2s and 2s (compared to 12s in Ethereum). The default order flow is First-Come First-Served (FCFS) with Priority Fee Auctions (PFA). On April 17, 2025, Arbitrum introduced the TimeBoost mechanism (explained further).
@@
Methodology. A key challenge in our empirical analysis is that DeFi smart contracts do not emit event logs for reverted transactions. To overcome this, we analyze transaction traces directly. For each reverted transaction, we construct a trace-based execution graph and search it to identify DEX or router interactions, targeted liquidity pools (Uniswap v2/v3), and token addresses (Uniswap v4). Although all analyzed rollups are EVM-compatible, each liquidity pool, AMM DEX, router, and token contract has a different address across L2s, requiring chain-specific mapping. To support this, we build a labeling library for each rollup using event logs from successfully executed swaps, which we then use to match and annotate nodes in the trace graphs of reverted transactions.
Successful transactions. We extract swap details from Uniswap v2, v3, and v4 by parsing their emitted event logs.
Reverted transactions. Since reverted transactions do not emit logs, we use on transaction traces to analyze their behavior.
@@
Bot Identification.
MEV bots are typically smart contracts storing the strategy and logic for interacting with DeFi protocols or other smart contracts. Since these bots must be externally triggered to execute MEV strategies, we identify potential bots by analyzing the to addresses in reverted transactions. Specifically, we focus on those addresses that are frequently targeted, excluding known DEX routers and AMM pools. To refine this classification, we apply the following criteria:
  • Addresses with a known label or owner name (e.g., wallets labeled by blockscan, centralized exchanges, or DAOs) are excluded, as they are unlikely to be bots. This is captured by the condition owa.owner_key IS NOT NULL.
  • Addresses that contain on-chain bytecode are flagged as likely automated agents (i.e., bots or smart contract traders), corresponding to the condition cm.contract_address IS NOT NULL.
@@ 5 Empirical Analysis
In the subsequent sections, we analyze the reverted transactions with respect to their use of priority fee auctions and their position within blocks. We classify reverted transactions into swap transactions and examine their key characteristics, including the targeted DEX and the token pair pool.
@@ 5.1 Priority-Fee Transactions
Further analysis of gas fees paid by reverted transactions is provided in Appendix 0.A, where we break down the total fee into three components: priority fee, base fee, and L1 data availability fee. Overall, although many reverted transactions either do not pay a priority fee or pay only a minimal one, the share of the priority fee in the total fee paid on L2s exceeds half across all analyzed L2 networks.
@@ 6 Discussion
The Dencun Effect. The Dencun upgrade introduced a significant shift in L2 gas economics. Pre-upgrade, the relatively high cost (often \$0.5–\$1 per transaction) disincentivized failed spam-based arbitrage. Post-Dencun, with average gas costs dropping below \$0.01, the marginal cost of failure became negligible. This cost asymmetry makes spam-based arbitrage economically rational post-upgrade and explains the observed spike in revert rates.
@@
Incentive Design and New Sequencing Mechanisms. The current status quo—FCFS ordering on centralized, private mempools—favors low-latency spamming and limits the scope for efficient, market-based prioritization. Several mechanisms have been proposed to address this:
@@
Underreliance on Priority Fee Auctions. A notable empirical observation is the limited use of priority fee auctions (PFA) by MEV searchers on rollups. Despite the theoretical promise of PFAs to prioritize transaction inclusion, we find that searchers often prefer to duplicate their transactions—submitting the same swap multiple times—rather than bid for priority.
@@ 7 Conclusions
Acknowledgments
The first author gratefully acknowledges support from the Flashbots Research Grant. The authors would also like to thank all collaborators from Flashbots and other institutions involved in the discussions that shaped this research.
@@ Appendix 0.A Gas Fees of Reverted Transactions
  • L2 Execution Fee. The cost of executing the transaction on the rollup’s virtual machine.
  • L1 Data Availability Fee. The cost of posting transaction calldata to Ethereum L1, which ensures data availability and security.